if A and B are acute angles and sin A = 1/5 √2 and tan B = 1/3 then A+2B=?

  1. if A and B are acute angles and sin A = 1/5√2 and tan B = 1/3 then A+2B=?

Grade:11

1 Answers

Akshay
8 years ago
 
Sin A= 1/5*rt(2), cos A=7/5*rt(2);        cos A= sqrt(1-sin2A)
Tan B= 1/3, sin B= 1/rt(10), cos B= 3/rt(10);  cos B= sqrt(1-sin2B)
sin (2B) = 2sinBcosB = 3/5, cos(2B) = 4/5;
Sin(A+2B) = sin(A)*cos(2B) + cos(A)*sin(2B);
= (4+21) /(25.rt(2))
=1/rt(2);
A+2B = 45 deg